OpenStudy (anonymous):

**MEDAL AND FAN** Raising any base to the power of zero will yield _____. A.1 B.0 C.–1 D.a nonreal number

zepdrix (zepdrix):

When you raise something to the 2nd power, you multiply it by itself. Example: 4^2 = 4*4 = 16 The exponent told us to multiply two sevens together. When you raise something to the 1st power, you don't change it. When you raise something to the 0th power, you divide the number out of itself. Example 4^0 = 4/4 = ?

OpenStudy (anonymous):

1 except 0^0 is undefined

zepdrix (zepdrix):

So what will that always give you?

OpenStudy (anonymous):

That gives me 1. Right?

OpenStudy (anonymous):

1

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

Okay thank u guys! Would u mind helping me in other questions pleasse????

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

Which of the following is a quantity that could be represented by the number 2.404x10^6 A.the population of Iowa B.the diameter (in meters) of an atom C.the thickness (in meters) of a sheet of printer paper D.the amount of money (in US dollars) in the average checking account

zepdrix (zepdrix):

Do you know how to write 2.404 x 10^6 in standard form?

OpenStudy (anonymous):

2.404000000 I think

zepdrix (zepdrix):

If you know how to convert it to standard form, it will help you see that it's a really really big number, which can only apply to one of our options.

zepdrix (zepdrix):

2,404,000 Big! :O Hmm so what could that describe? It's way too large to describe a sheet of paper in meters, right?

OpenStudy (anonymous):

Yes .

zepdrix (zepdrix):

Hmm so which one do you think? :) Does the average person have 2,404,000 dollars in their bank account?